OpenStudy (anonymous):

partial variation equation - w=-7t-20 , find value of t for following value of w.

OpenStudy (vishal_kothari):

for following value of w?????????

OpenStudy (anonymous):

it says A Partial Variation is given by the equation w=-7t-20, find value of t for the following value of w. w =1

OpenStudy (vishal_kothari):

t=3....

OpenStudy (anonymous):

how???

OpenStudy (vishal_kothari):

it's not 3...

OpenStudy (vishal_kothari):

w=7t+20 1=7t+20 -19=7t t=-19/7

OpenStudy (anonymous):

so we changed it to positive?

OpenStudy (vishal_kothari):

yes...